P 3-24 (similar to) Your brother has offered to give you either $20,000 today or $40,000 in 10 years. If the interest rate is 7% per year, which option is preferable? The present value of the future amount (amount received n 10 years) is $1 (Round to the nearest dollar.)Your bank pays 2.5% interest per year. You put $1,300 in the bank today and $700 more in the bank in one year. How much will you have in the bank in two years? In two years, the amount that you will have is SO (Round to the nearest cent.)P 3-20 (similar to) Consider the following alternatives: i. $150 received in one year ii. $230 received in five years ili. $320 received in 10 years a. Rank the alternatives from most valuable to least valuable if the interest rate is 12% per year. b, what is your ranking if the interest rate is 4% per year? C. What is your ranking if the interest rate is 17% per year? a. Rank the alternatives from most valuable to least valuable if the interest rate is 12% per year. First calculate the present value (PV) of each alternative: The PV of $150 received in one year if the interest rate is 12% per year is $ Round to the nearest cent.
Answer #1

Present value of future amount is calculated using the PV function in excel in the below quesitons:

1) =PV(7%,10,0,40000) = 20,334

40,000 in 10 years is more preferable.

2) 1300 will stay in the bank for 2 years, but 700 will earn interest only for one year, hence 2 present value will be calculated and added.  

=PV(2.5%,2,0,1300) = 1237.36

=PV(2.5%,1,0,700) = 682.93

Total = 1,920

3)

PV of 150 in one year at 12% =PV(12%,1,0,150) = 134

PV of 230 in 5 years at 12% =PV(12%,5,0,230) = 131

PV of 320 in 10 years at 12% =PV(12%,10,0,320) = 103